Oriole Pet supply uses the conventional retail method to determine its ending inventory at cost. Assume the beginning inventory at cost (retail) were $787800 ($971700), purchases during the current year at cost (retail) were $3025800 ($3978300), freight in on these purchases totaled $182 700 sales during the our ent year totaled $3876000 and net ar ups markdowns ere $5100 $279900). what is the ending inventory value at cost? Hint: Round intermediate calculation to 3 decimal places, ลู 0.635 and final answer to 0 decimal places O$799200. 866718. 644954. $1074000
Answer #1

ANSWER = c) $644954

Cost Retail
Beginning inventory 787800 971700
Purchases 3025800 3978300
Freight 182700
Merchandise available for sale (cost) 3996300 4950000
Add: Markups 5100
Retail with Markups 4955100
Subtract: Markdowns (279900)
Merchandise available for sale (retail) 4675200
Less: Sales 3876000
Ending Inventory at Retail 799200

To get the ending inventory at cost, first calculate the cost-to-retail ratio. With the conventional method you calculate the ratio using only markups.
= 3,996,300 / 4,955,100 = 0.807

Ending inventory at cost = 799200 * 0.807

= $ 644,954